the roots of the equation z^4+az^3+(12+9i)az^2+bz=0( where a nd b are complex numbers) are the vertices of a square


1
Expert's answer
2021-07-28T06:48:57-0400

Let "c" be the center of the square, and "r" be the distance from the center to vertices.

Then the roots of the given polynomial can be written as "c+re^{i\theta}, c+ire^{i\theta}, c-re^{i\theta}, c-ire^{i\theta}" for some "\theta\in[-\pi,\pi)". Therefore the given polynomial must be equal to the product

"(z-(c+re^{i\theta})) (z-(c+ire^{i\theta}))(z-(c-re^{i\theta}))(z-(c-ire^{i\theta}))="

"((z-c)-re^{i\theta})( (z-c)+re^{i\theta})((z-c)-ire^{i\theta})((z-c)+ire^{i\theta}))="

"((z-c)^2-r^2e^{2i\theta})((z-c)^2+r^2e^{2i\theta})=(z-c)^4-r^4e^{4i\theta}="

"z^4-4cz^3+6c^2z^2-4cz+c^4-r^4e^{4i\theta}"

Comparing the coefficients, we obtain

"a=b=-4c", "c^4=r^4e^{4i\theta}" and "6c^2=(12+9i)a=-4c(12+9i)".

From the last equation either c=0 or "c=-8-6i".

The case c=0 is degenerated, since in this case r=0 and the square is, in fact, a point.

We do not consider this case. Therefore, "c=-8-6i", "r=|c|=\sqrt{(-8)^2+(-6)^2}=10"

"e^{i\theta}=\frac{c}{r}e^{\frac{2\pi i}{4}k}=(-\frac{4}{5}-\frac{3}{5}i)e^{\frac{\pi ik}{2}}" with k=0,1,2 and 3 (each value of k is suitable).

"a=b=-4c=-4(-8-6i)=32+24i"

Answer. "a=b=32+24i".